# Rethinking Irregular Time Series Forecasting: A Simple yet Effective Baseline

### v1

@@ Abstract
The forecasting of irregular multivariate time series (IMTS) is crucial in key areas such as healthcare, biomechanics, climate science, and astronomy. However, achieving accurate and practical predictions is challenging due to two main factors. First, the inherent irregularity and data missingness in irregular time series make modeling difficult. Second, most existing methods are typically complex and resource-intensive. In this study, we propose a general framework called APN to address these challenges. Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching. By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner. Additionally, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency. Finally, predictions are made by a shallow MLP. Experimental results on multiple real-world datasets show that APN outperforms existing state-of-the-art methods in both efficiency and accuracy.
@@ 1 Introduction
Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao et al., 2018; Brouwer et al., 2019; Kidger et al., 2020; Ham et al., 2019; Vio et al., 2013). Irregular Multivariate Time Series Forecasting (IMTSF) is a crucial research task that provides valuable insights for early warning and proactive decision-making. However, the inherent irregularity of observations and data missingness (Yalavarthi et al., 2024; Zhang et al., 2024) in IMTS pose significant challenges to IMTSF modeling.
To address this challenge, recent IMTSF methods, such as T-PatchGNN (Zhang et al., 2024) and TimeCHEAT (Liu et al., 2025), have adopted the Fixed Patching approach—see Figure 1a. This approach divides the time series into fixed-length patches with equal intervals among them. However, this approach has notable limitations: 1) Uneven Information Density Across Patches: Fixed Patching struggles to adapt to local variations in data density, leading to uneven information density among patches. For example, sparse patches (where the number of observations is limited) may result in insufficient feature extraction, yet dense patches (where the number of observations is abundant) may contain redundant information or noise, thereby impairing the extracted features. 2) Inappropriate Segmentation of Key Semantic Information: Fixed Patching risks splitting critical dynamic information, which hampers the model’s ability to capture the complete semantic context. Therefore, the first challenge is: how to design an adaptive patching approach that can adapt to the local information density variations of IMTS and capture complete semantic information.
Meanwhile, existing IMTSF models generally suffer from high computational costs and long running times. For example, neural ODEs-based models (Chen et al., 2018; Rubanova et al., 2019; Gravina et al., 2024) require computationally intensive numerical solvers to accurately model continuous-time dynamics (Shukla and Marlin, 2021; Chen et al., 2018). GNNs-based models (Yalavarthi et al., 2024; Zhang et al., 2024) are hindered by the overhead of complex
@@
graph construction and multi-round node information aggregation (Wu et al., 2021); Transformer-based models (Chen et al., 2023; Zhang et al., 2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al., 2024). These models typically construct computationally intensive and parameter-heavy complex architectures to effectively handle the intricate dependencies and dynamic changes in IMTS. While this design enhances the model’s performance to some extent, it also incurs high computational costs and long runtimes, limiting its practical application in resource-constrained scenarios. Notably, in regular time series forecasting task, models such as DLinear (Zeng et al., 2023), SparseTSF (Lin et al., 2024a), and CycleNet (Lin et al., 2024b) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters. Therefore, the second challenge is: how to design an effective model for IMTSF.
To address the above challenges, we propose a general framework called APN. Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching—see Figure 1b. By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner. Based on the representations, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency. Finally, predictions are made by a shallow MLP. Results on multiple real-world datasets show that APN outperforms existing state-of-the-art methods in both efficiency and accuracy.
@@
  • To address the IMTSF, we propose a general framework named APN. This framework leverages adaptive patching to generate high-quality and regular initial patch representations. Based on these representations, we employ a simple query module to integrate contextual information, ensuring the effective design of the framework.
  • We design a novel TAPA module to achieve adaptive patching. By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequence into high-quality, regularized representations in a channel-independent manner, effectively adapting to local variations in information density and capturing complete semantics.
  • We conduct extensive experiments on multiple datasets. The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency.
@@ 2.2 Progress in Patch-based Irregular Multivariate Time Series Forecasting
Segmenting time series into “patches” to capture local patterns has demonstrated remarkable success in regular time series forecasting tasks, as evidenced by models like PatchTST (Nie et al., 2023) and Crossformer (Wang et al., 2022). Inspired by this, the patch-based approach has also been applied to IMTSF. T-PatchGNN (Zhang et al., 2024) employs its “Transformable Patch” mechanism to convert irregular sequences into patches with uniform temporal resolution but variable lengths, thereby maintaining temporal alignment while adapting to the sparsity and density variations of observation points. Meanwhile, TimeCHEAT (Liu et al., 2025) divides IMTS into fixed-length patches and utilizes graph learning within each patch for fine-grained temporal embedding. Both methods ingeniously leverage the patch strategy to transform raw IMTS data into structured representations that are more amenable to model processing, laying the groundwork for subsequent sequence modeling.
Unlike the fixed patching methods mentioned above, this study introduces adaptive patching. By learning dynamically adjustable patch boundaries and employing a time-aware weighted averaging strategy, our method transforms the original irregular sequence into a high-quality regularized patch representation in a channel-independent manner. This approach effectively adapts to local variations in information density and captures complete semantic information.
@@ 3.2 Framework Overview
The overall pipeline of APN is shown in Figure 2. Its core component is the Time-Aware Patch Aggregation (TAPA) Module–see Section 3.3, which processes the multivariate irregular sequence in a Channel-Independent manner. Compared with conventional fixed segmentation, TAPA utilizes the Adaptive Patching and Weighted Aggregation. This design explicitly aims to adapt to the local variations in information density inherent in IMTS and to capture complete semantic information. Then, a concise and efficient Query-based Aggregation module–see Section 3.4 generates a representation of historical information for each variable. Finally, this representation is fed into a simple MLP-based Forecasting decoder to make predictions–see Section 3.5. In brief, the core advantage of APN stems from the high-quality representations produced by TAPA, which effectively compresses the historical information through adapting to variations in the distribution and concentration of information and preserving complete semantic context, thus APN achieves state-of-the-art performance with a subsequent simple MLP decoder.
@@ 3.3 Time-Aware Patch Aggregation (TAPA)
Traditional Fixed-Span Patching (Zhang et al., 2024; Liu et al., 2025), due to the fixed patch boundaries or predefined time spans, struggles to adapt to local variations in the underlying information content and its temporal distribution (referred to as information density), and to capture complete semantic information in IMTS. To overcome this limitation, the TAPA mechanism devises the Adaptive Patching and Weighted Aggregation to (1) adapt the patching strategy to the local information density variations within each time series, and (2) ensure that crucial semantic information is holistically captured within the generated patches by focusing on the information most important for the prediction task. It efficiently integrates the information in each univariate irregular sequence $X^{n}=\{(t^{n}_{i},v^{n}_{i})\}_{i=1}^{L_{n}}$ by transforming it into regularized patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$. This process is performed independently for each variable, accommodating their unique sampling patterns and information distributions.
3.3.1 Adaptive Patching
The Adaptive Patching works through dynamically learning the temporal coverage range $[t_{p}^{left,n},t_{p}^{right,n}]$ for each target patch. This dynamic learning of boundaries is the primary mechanism by which TAPA adapts to local information density variations. For instance, in sparsely observed regions that nonetheless contain critical information, the learned patch width (controlled by $\lambda_{p}^{n}$) might expand to ensure this information is captured, or its position (controlled by $\delta_{p}^{n}$) might shift to center on these information-rich sparse points. Conversely, in densely observed regions with redundant information, a patch might become narrower to focus only on the novel or most salient signals, thus avoiding dilution by less informative data. Furthermore, this adaptive process enables the model to determine the content of each patch based on the inherent importance of the data, which is crucial for capturing complete semantic information. Specifically, TAPA learns the left boundary offset $\delta_{p}^{n}$ and a log-scale parameter $\lambda_{p}^{n}$ representing the patch width:
@@ 3.3.1 Adaptive Patching
where $\delta_{p}^{n}$ and $\lambda_{p}^{n}$ are learnable parameters, $c_{p}=(p-0.5)\cdot(T_{obs}/P)$ and $S_{init}=T_{obs}/P$ denote the reference center points and initial average patch width. For observations highly important for prediction (i.e., regions with high information density relevant to the task), different patch boundaries are adjusted to accurately cover these crucial regions, ensuring their semantic integrity. While for regions with noise or low information relevance (low information density), boundaries may contract or shift to reduce their coverage, thus preserving the clarity of the extracted semantic information. The dynamic adjustment of boundaries thus constitutes an effective mechanism for adaptively selecting useful information and forming semantically coherent patches based on learned information importance.
3.3.2 Weighted Aggregation
@@
To fully utilize temporal information from observations, the Learnable Time Embedding $TE(t_{i})\in\mathbb{R}^{D_{te}}$ is first generated for $i$-th timestamp, which consists of a linear layer for scale and a sine-activated linear layer for periodicity. Then the Learnable Time Embedding is concatenated with the original observed value $v_{i}$ to form an augmented representation $\tilde{v}_{i}=\text{Concat}(v_{i},TE(t_{i}))$.
After determining the dynamic coverage range $[t_{p}^{left,n},t_{p}^{right,n}]$ (which has already adapted to local information densities and aimed to cover semantic regions), we employ a query-like Weighted Aggregation. This weighting mechanism synergizes with the adaptive boundaries to further refine the captured semantic information by emphasizing the most salient points within each adaptively formed patch. For any observation $i$ (timestamp $t_{i}$) of variable $n$, its relationship $\alpha_{i,p}^{n}$ to target patch $p$ is defined by a soft window function:
@@
$\alpha_{i,p}^{n}=\sigma\left(\frac{t_{p}^{right,n}-t_{i}}{\text{Softplus}(\kappa^{n})}\right)\cdot\sigma\left(\frac{t_{i}-t_{p}^{left,n}}{\text{Softplus}(\kappa^{n})}\right),$ (4)
where $\sigma(\cdot)$ is the Sigmoid function, and $\text{Softplus}(\kappa^{n})$ is a temperature scaling factor controlled by a learnable scalar parameter $\kappa^{n}$, regulating the smoothness of the soft window boundary transition. Through learning the $\kappa^{n}$, the observations within the dynamically determined patch boundaries are further weighted, where important observations contributing to high local information density receive larger $\alpha_{i,p}^{n}$ in aggregation, while less relevant observations are down-weighted or filtered out, thus ensuring the completeness and purity of the captured semantic information.
After calculating the relationships between observations and target patches, the aggregated representation $\bar{h}_{p}^{n}$ for the $p$-th patch of variable $n$ can be obtained through a weighted average of the augmented features $\tilde{v}^{n}_{i}$:
$\bar{h}_{p}^{n}=\frac{\sum_{i}\alpha_{i,p}^{n}\cdot\tilde{v}^{n}_{i}}{\sum_{i}\alpha_{i,p}^{n}+\epsilon}\in\mathbb{R}^{1+D_{te}},$ (5)
where $\epsilon$ is a small positive constant to prevent division by zero. Subsequently, we transform $\bar{h}_{p}^{n}\in\mathbb{R}^{(1+D_{te})}$ into the uniform hidden representations $h_{p}^{n}\in\mathbb{R}^{D}$ to enhance the expressiveness through a linear projection layer $h_{p}^{n}=\text{Linear}_{D}(\bar{h}_{p}^{n})$.
Through the aforementioned mechanisms, each univariate irregular sequence $X^{n}$ is efficiently transformed into a refined and structurally regularized patch representation sequence $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]\in\mathbb{R}^{P\times D}$.

### v2

@@ Abstract
The forecasting of irregular multivariate time series (IMTS) is a critical task in domains like healthcare and climate science. However, this task faces two significant hurdles: 1) the inherent non-uniformity and missing data in IMTS complicate the modeling of temporal dynamics, and 2) existing methods often rely on computationally expensive architectures. To address these dual challenges, we introduce APN, a general and efficient forecasting framework. At the core of APN is a novel Time-Aware Patch Aggregation (TAPA) module that introduces an aggregation-based paradigm for adaptive patching, moving beyond the limitations of fixed-span segmentation and interpolation-based methods. TAPA first learns dynamic temporal boundaries to define data-driven segments. Crucially, instead of resampling or interpolating, it directly computes patch representations via a time-aware weighted aggregation of all raw observations, where weights are determined by each observation’s temporal relevance to the segment. This approach provides two key advantages: it preserves data fidelity by avoiding the introduction of artificial data points and ensures complete information coverage by design.The resulting regularized and information-rich patch representations enable the use of a lightweight query module for historical context aggregation and a simple MLP for final prediction. Extensive experiments on multiple real-world datasets demonstrate that APN establishes a new state-of-the-art, significantly outperforming existing methods in both prediction accuracy and computational efficiency.
@@ 1 Introduction
Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao et al., 2018; Brouwer et al., 2019; Kidger et al., 2020; Ham et al., 2019; Vio et al., 2013). Irregular Multivariate Time Series Forecasting (IMTSF) is a crucial research task that provides valuable insights for early warning and proactive decision-making. However, the inherent irregularity of observations and missing data (Yalavarthi et al., 2024; Zhang et al., 2024) in IMTS pose significant challenges to IMTSF modeling.
To address this challenge, recent IMTSF methods, such as T-PatchGNN (Zhang et al., 2024) and TimeCHEAT (Liu et al., 2025), have adopted the Fixed Patching approach—see Figure 1a. This approach divides the time series into fixed-length patches with equal intervals among them. However, this approach has notable limitations: 1) Uneven Information Density Across Patches: Fixed Patching struggles to adapt to local variations in data density, leading to uneven information density among patches. For example, sparse patches (where the number of observations is limited) may result in insufficient feature extraction, yet dense patches (where the number of observations is abundant) may contain redundant information or noise, thereby impairing the extracted features. 2) Inappropriate Segmentation of Key Semantic Information: Fixed Patching risks splitting critical dynamic information, which hampers the model’s ability to capture the complete semantic context. Therefore, the first challenge is to design an adaptive patching approach that can effectively adapt to local information density variations in IMTS and capture the complete semantic context.
@@
Meanwhile, existing IMTSF models generally suffer from high computational costs and long running times. For example, neural ODEs-based models (Chen et al., 2018; Rubanova et al., 2019; Gravina et al., 2024) require computationally intensive numerical solvers to accurately model continuous-time dynamics (Shukla and Marlin, 2021; Chen et al., 2018). GNNs-based models (Yalavarthi et al., 2024; Zhang et al., 2024) are hindered by the overhead of complex graph construction and multi-round node information aggregation (Wu et al., 2021); Transformer-based models (Chen et al., 2023; Zhang et al., 2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al., 2024). These models typically construct computationally intensive and parameter-heavy complex architectures to effectively handle the intricate dependencies and dynamic changes in IMTS. While this design enhances the model’s performance to some extent, it also incurs high computational costs and long runtimes, limiting its practical application in resource-constrained scenarios. Notably, in regular time series forecasting task, models such as DLinear (Zeng et al., 2023), SparseTSF (Lin et al., 2024b), and CycleNet (Lin et al., 2024a) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters. Therefore, the second challenge is to design a model for IMTSF that is both highly accurate and computationally efficient.
To address the above challenges, we propose a general framework called APN. Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching—see Figure 1b. Unlike methods that resample data points within a patch, TAPA directly aggregates the raw, irregular observations. It learns dynamically adjustable patch boundaries to flexibly frame regions of interest. Crucially, it then employs a time-aware soft-weighting mechanism to generate a patch representation, ensuring that all observations are holistically considered and contribute to the representation based on their temporal relevance to the patch. This approach not only adapts to local information density and captures complete semantic units but also inherently prevents information loss by design, as no data point is left uncovered. Based on these high-quality representations, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency. Finally, predictions are made by a shallow MLP. Results on multiple real-world datasets show that APN outperforms existing state-of-the-art methods in both efficiency and accuracy.
@@
  • To address the IMTSF, we propose a general framework named APN. This framework leverages adaptive patching to generate high-quality and regular initial patch representations. Based on these representations, we employ a simple query module to integrate contextual information, enabling a lightweight yet powerful overall architecture.
  • We design a novel TAPA module to achieve adaptive patching. By learning dynamic patch boundaries and employing a time-aware soft-aggregation scheme over raw observations, TAPA produces high-quality regularized representations without interpolation. This mechanism effectively adapts to local variations in information density, captures complete semantics, and guarantees full coverage of all data points.
  • We conduct extensive experiments on multiple datasets. The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency. Additionally, all datasets and code are available at https://anonymous.4open.science/r/APN-23D0.
@@ 2.2 Progress in Patch-based Irregular Multivariate Time Series Forecasting
The patch-based strategy, which has proven successful for regular time series (Nie et al., 2023; Wang et al., 2022), has been adapted for IMTS Forecasting (IMTSF). Initial adaptations, such as T-PatchGNN (Zhang et al., 2024) and TimeCHEAT (Liu et al., 2025), employed a fixed-span patching strategy. However, this rigid segmentation is ill-suited for the non-uniform data distribution of IMTS. It often creates patches with highly variable information content—some being too sparse for meaningful feature extraction, while others may be overly dense with redundant data.
To overcome this, adaptive patching emerged as a natural evolution. One prominent approach, primarily explored in the context of regular time series (e.g., HDMixer (Huang et al., 2024)), involves learning flexible patch boundaries and then generating regularized patch representations through interpolation-based resampling. While effective for dense, regular data, this methodology is fundamentally flawed for IMTS. Interpolating across sparse or missing intervals can fabricate misleading data points, introducing significant artifacts and undermining model reliability.
In contrast, our work proposes a distinct aggregation-based adaptive patching paradigm, specifically designed for the challenges of IMTS. Instead of creating new, artificial data points, our TAPA module learns dynamic "soft windows" and computes each patch representation by performing a direct, weighted aggregation of all original observations. This strategy provides two critical advantages for IMTS:(1)Data Fidelity: It exclusively uses the original, observed data, thereby avoiding the distortions and potential inaccuracies of interpolation.(2)Complete Information Coverage: The soft-weighting mechanism ensures that every data point contributes to the representation of relevant patches, mathematically precluding the information loss that can occur when observations fall between the hard boundaries of other methods.
@@ 3.2 Framework Overview
The architectural design of APN is guided by a core principle: decoupling the challenge of handling irregularity from the task of forecasting. As illustrated in Figure 2, instead of relying on a monolithic, complex model, APN adopts a strategic two-stage pipeline.
The cornerstone of this pipeline is the Time-Aware Patch Aggregation (TAPA) module (Section 3.3). Operating in a channel-independent manner, its sole yet critical purpose is to transform the raw, irregular observations into a high-quality, regularized sequence of patch representations. Through its novel use of Adaptive Patching and Weighted Aggregation, TAPA robustly adapts to local information density and captures complete semantic units. Crucially, it achieves this without resorting to interpolation, thereby preserving data fidelity and sidestepping the introduction of artificial artifacts.
The success of this initial transformation is pivotal, as it enables the subsequent use of a remarkably efficient and lightweight architecture. A concise Query-based Aggregation module (Section 3.4) then effectively summarizes the historical context from these regularized patches for each variable. Finally, this compact representation is fed into a simple MLP-based Forecasting Decoder (Section 3.5) for the final prediction.
In essence, APN strategically front-loads the complexity of handling irregularity into the novel TAPA module. By producing information-rich, regularized representations upfront, it obviates the need for computationally expensive back-end models, achieving state-of-the-art performance through an elegant and efficient design.
@@ 3.3 Time-Aware Patch Aggregation (TAPA)
At the heart of APN lies the Time-Aware Patch Aggregation (TAPA) module, a novel mechanism designed to fundamentally reframe how we process irregular time series. Traditional methods rely on a paradigm of "hard segmentation," imposing rigid, fixed-span boundaries that are ill-suited to the non-uniform nature of IMTS. In stark contrast, TAPA introduces a "soft aggregation" paradigm. Instead of discretely assigning observations to patches, it conceptualizes each patch as a soft, overlapping field of influence, allowing it to aggregate information from all observations in a temporally-aware manner.
This paradigm shift is realized through a cohesive, two-step process executed independently for each channel $X^{n}=\{(t^{n}_{i},v^{n}_{i})\}_{i=1}^{L_{n}}$. First, in Adaptive Patching, the model learns the dynamic temporal characteristics—the center and scale—of each patch’s field of influence. Second, in Weighted Aggregation, it computes a representation for each patch by performing a direct, weighted aggregation of all raw observations, where the weights are determined by each observation’s temporal relevance to the patch’s learned characteristics. This entire process transforms the irregular input sequence into a regularized, information-rich sequence of patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$.
3.3.1 Adaptive Patching: Learning Dynamic Temporal Windows
The first step is to define the temporal scope of each patch. Instead of imposing pre-defined, rigid boundaries, TAPA learns a dynamic temporal window, $[t_{p}^{left,n},t_{p}^{right,n}]$, for each target patch $p$. This dynamism is the core mechanism by which TAPA adapts to the non-uniform distribution of data. For instance, in a sparse but critical region, the model can learn to expand a window’s width to ensure sparse yet vital signals are encompassed. Conversely, in a dense region with redundant data, a window can be strategically narrowed to focus on the most salient signals, thus preventing dilution from less informative points.
This adaptive process empowers the model to carve out semantically coherent segments from the raw time series, based on the data’s inherent structure. To achieve this, for each patch $p$ and channel $n$, TAPA learns two key parameters: a positional adjustment $\delta_{p}^{n}$ and a log-scale width parameter $\lambda_{p}^{n}$. The window boundaries are then computed as:
@@ 3.3.1 Adaptive Patching: Learning Dynamic Temporal Windows
where $T_{obs}$ is the time span of the historical observation window, and $c_{p}=(p-0.5)\cdot(T_{obs}/P)$ and $S_{init}=T_{obs}/P$ represent the initial reference center and width, respectively. The use of $\exp(\lambda_{p}^{n})$ is a deliberate design choice, ensuring the learned window width remains strictly positive throughout optimization. This dynamic windowing mechanism allows the model to adaptively frame regions of interest, laying the foundation for a more meaningful and robust aggregation in the subsequent step.
3.3.2 Weighted Aggregation: From Raw Observations to Rich Representations
With the dynamic temporal windows defined, the next step is to generate patch representations by aggregating information from the raw observations.
@@ 3.3.2 Weighted Aggregation: From Raw Observations to Rich Representations
Enriching Observations with Temporal Context. To fully leverage temporal information, we first enrich each raw observation. A learnable time embedding, $TE(t_{i})\in\mathbb{R}^{D_{te}}$, is generated for each timestamp $t_{i}$. This embedding, composed of linear layers for scale and sine-activated layers for periodicity, captures both the absolute position and periodic patterns in time. It is then concatenated with the original value $v_{i}$ to form an augmented representation: $\tilde{v}_{i}=\text{Concat}(v_{i},TE(t_{i}))$.
The Soft Window Function. We then employ a time-aware soft window function to calculate the relevance weight, $\alpha_{i,p}^{n}$, of each observation $i$ to each patch $p$. The function is elegantly constructed as the product of two Sigmoid functions:
@@
$\alpha_{i,p}^{n}=\sigma\left(\frac{t_{p}^{right,n}-t_{i}}{\text{Softplus}(\kappa^{n})}\right)\cdot\sigma\left(\frac{t_{i}-t_{p}^{left,n}}{\text{Softplus}(\kappa^{n})}\right).$ (4)
Here, the first Sigmoid term models a smooth "falloff" from the right boundary, while the second models a smooth "rise" from the left boundary. Their product creates a continuous, bell-shaped weighting curve centered within the patch’s learned range. The term $\text{Softplus}(\kappa^{n})$ is a learnable, strictly positive temperature that controls the softness of the window edges. A smaller temperature leads to sharper, more defined boundaries, while a larger temperature creates gentler, more overlapping fields of influence, granting the model further adaptive flexibility.
A Crucial Design Choice for Information Integrity. The choice of this soft window function is critical. Since the Sigmoid function $\sigma(x)$ is strictly positive for all real inputs, the resulting weight $\alpha_{i,p}^{n}$ is also guaranteed to be positive for any observation $i$ and any patch $p$. This property is not a mere side effect; it is a fundamental design principle that mathematically guarantees complete information coverage. Every observation contributes to the representation of every patch (albeit with varying degrees of influence), inherently preventing the information loss that plagues all hard-segmentation methods where observations can be inadvertently discarded.
Final Aggregation. The final representation for the $p$-th patch, $\bar{h}_{p}^{n}$, is computed as a normalized weighted average of all augmented observation features:
$\bar{h}_{p}^{n}=\frac{\sum_{i=1}^{L_{n}}\alpha_{i,p}^{n}\cdot\tilde{v}^{n}_{i}}{\sum_{i=1}^{L_{n}}\alpha_{i,p}^{n}+\epsilon}\in\mathbb{R}^{1+D_{te}},$ (5)
where the summation spans all $L_{n}$ observations, and $\epsilon$ ensures numerical stability. Finally, to enhance expressiveness and align dimensions, we project $\bar{h}_{p}^{n}$ into the model’s uniform hidden space via a linear layer: $h_{p}^{n}=\text{Linear}_{D}(\bar{h}_{p}^{n})$. Through this entire process, each univariate irregular sequence is transformed into a refined, structurally regularized sequence $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]\in\mathbb{R}^{P\times D}$.